OpenStudy (anonymous):

exact value of: tan112.5

OpenStudy (raden):

112.5 = 225/2 so, tan(112.5) = tan(225/2) use the half angle identity : tan(x/2) = sqrt ([1-cos(x)]/[1+cos(x)])

OpenStudy (anonymous):

thank you!

OpenStudy (raden):

the other formulas can u look here : http://www.mathwords.com/h/half_angle_identities.htm welcome :)